# Milky Way Globular Clusters: Nurseries for Dynamically-Formed Binary Black Holes

### v1

@@
11institutetext: Dipartimento di Fisica, Sapienza, Università di Roma, Piazzale Aldo Moro 5, 00185, Roma, Italy22institutetext: Dipartimento di Fisica, Tor Vergata, Università di Roma, via della Ricerca Scientifica, 00133, Roma, Italy33institutetext: INAF/Osservatorio Astronomico di Roma, Via di Frascati 33, 00078 Monte Porzio Catone, Italy44institutetext: INFN, Sezione di Roma I, Piazzale Aldo Moro 2, 00185 Roma, Italy55institutetext: William H. Miller III Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles Street, Baltimore, Maryland, 21218, USA66institutetext: Max Planck Institute for Astrophysics, Karl-Schwarzschild-Str. 1, 85748 Garching, Germany77institutetext: Physics and Astronomy Department Galileo Galilei, University of Padova, Vicolo dell’Osservatorio 3, I–35122, Padova, Italy88institutetext: INFN - Padova, Via Marzolo 8, I–35131 Padova, Italy99institutetext: Institut für Theoretische Astrophysik, ZAH, Universität Heidelberg, Albert-Ueberle-Straße 2, D-69120, Heidelberg, Germany1010institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, 69117, Heidelberg, Germany
Milky Way Globular Clusters: Nurseries for Dynamically-Formed Binary Black Holes
@@
(Received June 2025; accepted …)
We present a novel self-consistent theoretical framework to characterize the formation, evolution, and merger sites of dynamically-formed black hole binaries, with a focus on explaining the most massive events observed by the LIGO-Virgo-KAGRA Collaboration. Our approach couples the galaxy formation model GAMESH with cluster population synthesis codes to trace the cosmic evolution of globular clusters simultaneously with mergers of massive black holes.
Our reference model, which includes prescriptions for both cluster formation and disruption depending on properties of specific galaxies, accurately reproduces the observed age-mass distribution of the Milky Way’s globular clusters. We find that approximately 30% of the globular clusters observed in our galaxy’s halo may have originated from satellite galaxies of the Milky Way. We confirm that hierarchical black hole mergers provide a significant contribution to the formation of black holes in and above the pair-instability mass gap. However, quantifying their contribution is challenging, as different population synthesis codes yield divergent results in terms of black hole mass function and merger rates.
Furthermore, we characterize the host galaxies where massive black holes form in terms of their dark matter, stellar mass, and metallicity.
Ultimately, we demonstrate that the merger and birth rate densities of binary black holes increase with redshift till z = 5. This cosmic evolution is a crucial signature with significant implications for future detectors like the LISA, the Einstein Telescope and Cosmic Explorer, which will be capable to probe the high-redshift Universe.
@@ 1 Introduction
The detection of gravitational waves (GWs) in 2015 (Abbott+2016) marked the beginning of a new era in astrophysics, unveiling the mysterious universe of black holes (BHs). Since then, the LIGO-VIRGO-KAGRA (LVK) Collaboration has published nearly 180 binary BH (BBHs) (Abbott+2023; GWTC4v1+2025), providing crucial insights into the fundamental processes that govern the formation and evolution of these enigmatic compact objects (Abbott+2017; Abbott+2019; Abbott+2021; Abbott+2023BHPop; Abbott+2024; GWTC4+2025). Despite remarkable progresses, the properties of the merging BBHs detected by LVK, such as their mass and spin distributions, continue to challenge theoretical models. For example, the absence of a high-mass cutoff in the primary BH mass distribution above $\sim 60$ M (Farmer+2019; Farmer+2020; VanSon+2020; Marchant+2020; Callister+2024) has raised questions about the effective existence of a pair-instability mass gap. This gap, which is predicted by stellar evolution models, does not appear in the LVK catalog (Abbott+2023BHPop; GWTC4+2025), which includes several detections of BHs within the gap region (Abbott+2019). Nevertheless, the precise astrophysical channel responsible for the formation of massive stellar BBHs (MBBHs) with $_{\rm BH}>50 M$ remains elusive, with a variety of mechanisms proposed to explain the most massive GW events. These mechanisms include isolated binary evolution (Bethe+1998; Belczynski+2008; Dominik+2012; Dominik+2013; Dominik+2015; Belczynski+2016; Schneider+2017; Giacobbo+2018; Marassi+2019; Graziani+2020b; Broekgaarden+2022; Franciolini+2024), direct star collisions (DiCarlo+2019; Renzo+2020; Kremer+2020a; Kremer+2020b; Costa+2022; Ballone+2023), three-body interactions (Antonini+2017; ArcaSedda+2021; VignaGomez+2021), and dynamical formation channels involving star clusters (Mapelli+2016; Kumamoto+2020) and the disks of active galactic nuclei (McKernan+2012; Bartos+2017; McKernan+2018; Tagawa+2019; Tagawa+2020; Ishibashi+2020; Vaccaro+2024; McKernan+2024). Among the proposed mechanisms, dynamical formation in dense stellar environments - such as globular clusters (GCs) and nuclear star clusters - has attracted significant interest over the past few decades. These environments are particularly well-suited for the formation of hierarchical BH merger chains, where the compact remnants of massive stars can undergo successive coalescences, leading to the formation of increasingly massive BHs (Miller+2002; Holley-Bockelmann+2008; Giersz+2015; Antonini+2015; Antonini+2016; Mapelli+2021; Mapelli+2022; Chattopadhyay+2023; Kritos:2022non; Torniamenti+2024). Indeed, due to mass segregation, BHs gradually migrate toward the cluster core on sub-gigayear timescales, resulting in the formation of a BH-dominated central region (Spitzer+1969; Kulkarni+1993; Sigurdsson+1993). In such BH-dominated cores, dynamically hard BBHs are promptly formed through three-body interactions, and further hardened by GW capture, binary-single, and binary-binary scattering events (Phinney+1991; Portegies+2000; Peuten+2016; ArcaSedda+2018; Zocchi+2019; Kremer+2020a; Antonini+2020). Additionally, dynamical mechanisms can produce unique features which could help to distinguish the contribution of this formation channel from that of isolated binary evolution, such as large spin magnitudes inherited from the pre-merger orbital angular momentum (Gerosa+2017) and non-negligible eccentricities (Nishizawa+2016; Breivik+2016; Nishizawa+2017; Samsing+2019; Martinez+2020; Rom+2024). However, the dynamical formation of BBHs presents several challenges. A major obstacle is the relativistic kick imparted to merger remnants at birth (Campanelli+2007; Lousto+2011), which can exceed the escape velocity of their host clusters. This often results in the ejection of the remnants from their environment, thereby disrupting the hierarchical merger chain (Favata+2004; Holley-Bockelmann+2008; Kesden+2010). The interplay between the properties of the host environment and the strength of the kicks plays a critical role in shaping the mass distribution of the BH population resulting from this formation pathway (Islam:2025drw). Unfortunately, due to the large distances involved and limited sky localization accuracy, confidently determining the orbital properties and specific birth environments of BBHs remains challenging. This uncertainty hampers our ability to fully understand the true origin of MBBHs and to rule out alternative formation channels, such as binary evolution of Population III stars (Kinugawa+2014; Kinugawa+2016; Tanikawa+2021b; Wang+2022; Tanikawa+2022; Costa+2023; Santoliquido+2023; Mestichelli+2024; Liu+2024) and primordial BHs (DeLuca+2021; Ng:2022agi; ChenZu+2024; DeLuca+2025) as contributors to the observed GW events. As the number of detected mergers continues to grow with ongoing and future runs, the prospect of disentangling the contributions from different formation mechanisms becomes increasingly attainable. Bayesian inference methods (Thrane+2019; Abbott+2023BHPop; Franciolini+2022; Franciolini+2024) and population synthesis models (Hurley+2000; Portegies+1995; Mapelli+2013; Spera+2015; GiacobboMapelli+2018; Elbert+2018; Breivik+2020; Riley+2022; Iorio+2023) enable the inference of the intrinsic properties of observed BBHs by comparing observational data with predictions derived from various astrophysical scenarios and binary initial conditions. These statistical comparisons have opened new avenues for uncovering the true formation history behind these mergers, despite the significant uncertainties associated with stellar and binary evolution (Iorio+2023). In this work, we focus on the role of the dynamical formation scenario in dense stellar environments, with particular emphasis on the contribution of BBHs formed in GCs to the observed population of BH mergers. Additionally, we investigate how the distinctive properties of BHs resulting from hierarchical mergers may provide new insights into the astrophysical processes underlying the formation of MBBHs. To investigate this, various approaches have been employed. Traditionally, the numerical complexity of direct $N$-body and Monte Carlo simulations has limited studies of cluster dynamics to a relatively small number of cases (Spurzem+1999; Rodriguez+2015; Rodriguez+2016a; Rodriguez+2016b; Wang+2015; Wang+2016; Chatterjee+2017; Askar+2017; Rodriguez+2019; Banerjee+2020; ArcaSedda+2023; ArcaSedda+2024). However, the recent development of fast semi-analytical codes, also known as cluster population synthesis (CPS) codes – such as B-POP (ArcaSedda+2021), cBHBd (Antonini+2020; Antonini+2023), FASTCLUSTER (Mapelli+2021), QLUSTER (Gerosa+2023) and RAPSTER (Kritos+2024) – has made it possible to conduct cluster population studies within a reasonable time-frame. By combining CPS predictions with models of the cosmic cluster formation rate (Mapelli+2022; GonzalezPrieto+2022; Kritos+2024; Torniamenti+2024; Mestichelli+2024), one can rapidly estimate the average merger rate density of dynamically-formed BBHs as a function of redshift. However, this approach does not provide detailed information about the spatial distribution or the chemical properties of the galaxies hosting globular clusters. Understanding the properties of host galaxies is crucial for improving our knowledge of GCs and, by extension, of MBBH formation (see, e.g., Krumholz+2019 for a review). Several key aspects of cluster formation are known to vary with the galactic environment, including the cluster formation efficiency (Ginsburg+2018), the potential high-mass truncation of the cluster initial mass function (Wainer+2022), and the mass–radius relation (Brown+2021). Moreover, the time evolution of individual star clusters is strongly influenced by a range of environmental factors (e.g., Rossi+2016; Suin+2022). Therefore, a realistic model of GC formation and evolution must incorporate the evolving, galaxy-specific environmental conditions, and do so across a statistically significant sample of galaxies to adequately represent their diversity. By incorporating the progression of chemical enrichment in galaxies throughout their assembly, hydrodynamic galaxy formation simulations offer a powerful tool to study the population of dynamically-formed BBHs emerging from dense stellar environments. However, due to their high computational cost, these simulations require a trade-off between mass and spatial resolution, the size of the simulated volume, and the redshift range covered. Currently, most high-resolution zoom-in simulations – those with sub-parsec resolution capable of capturing star cluster formation – are limited to small cosmological volumes and/or high redshifts (e.g., Boley+2009; Kimm+2016; Lahen+2019; Calura+2022; Calura+2024). When the mass resolution is insufficient to resolve star cluster formation within galaxies, semi-analytical models become essential to complement these simulations. Studies utilizing this approach have already been carried out, demonstrating promising agreement with the observed GC populations in nearby galaxies (e.g., Li+2014; Pfeffer+2018). This study aims to investigate the physical properties of dynamically-formed BBHs originating in GCs during the assembly of the Local Group (LG) simulated by the galaxy formation model GAMESH (Graziani+2017). By estimating their merger rate density and examining their birth environments, we quantify the contribution of this dynamical formation channel to the GW events detected by the LVK Collaboration. Additionally, our analysis enables the identification of the host galaxies of these BBHs and the characterization of their physical properties. To this end, in Section 2 we outline the adopted galaxy formation model (2.1), in Section 2.2 we detail the coupling between the CPS framework and the galaxy evolution simulation, and in Section 2.3 we describe our cluster evolution models. Our main findings are presented in Section 3, followed by an extended comparison with previous studies in Section 4. This discussion is especially timely because of the recent release of O4 data (LIGOScientific:2025slb).
@@ 2 Methods
In this study, we adopt a LG-like volume simulation, named GAMESH (Graziani+2015) (Section 2.1) and propose a novel semi-analytic coupling approach for GC formation (Section 2.2), enabling the placement of cluster populations within each galaxy. Finally, we integrate the time evolution of GCs using the RAPSTER and FASTCLUSTER codes (Section 2.3), marking their first application within the context of a numerical galaxy formation simulation.

### v2

# Milky Way globular clusters: Nurseries for dynamically formed binary black holes

@@
11institutetext: Dipartimento di Fisica, Sapienza, Università di Roma, Piazzale Aldo Moro 5, 00185, Roma, Italy22institutetext: Dipartimento di Fisica, Tor Vergata, Università di Roma, via della Ricerca Scientifica, 00133, Roma, Italy33institutetext: INAF/Osservatorio Astronomico di Roma, Via di Frascati 33, 00078 Monte Porzio Catone, Italy44institutetext: INFN, Sezione di Roma I, Piazzale Aldo Moro 2, 00185 Roma, Italy55institutetext: William H. Miller III Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles Street, Baltimore, Maryland, 21218, USA66institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, 69117, Heidelberg, Germany77institutetext: Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany88institutetext: Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany99institutetext: Physics and Astronomy Department Galileo Galilei, University of Padova, Vicolo dell’Osservatorio 3, 35122 Padova, Italy1010institutetext: INFN – Padova, Via Marzolo 8, 35131 Padova, Italy1111institutetext: INAF – Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, 35122 Padova, Italy
Milky Way globular clusters: Nurseries for dynamically formed binary black holes
@@
(Received …; accepted …)
We present a novel self-consistent theoretical framework to characterize the formation, evolution, and merger sites of dynamically formed black hole binaries, with a focus on explaining the most massive events observed by the LIGO-Virgo-KAGRA Collaboration. Our approach couples the galaxy formation model GAMESH with cluster population synthesis codes to trace the cosmic evolution of globular clusters simultaneously with mergers of massive black holes.
Our reference model, which includes prescriptions for both cluster formation and disruption depending on properties of specific galaxies, accurately reproduces the observed age-mass distribution of the Milky Way’s globular clusters. We find that approximately 30% of the globular clusters observed in our galaxy’s halo may have originated from satellite galaxies of the Milky Way. We confirm that hierarchical black hole mergers provide a significant contribution to the formation of black holes in and above the pair-instability mass gap. However, quantifying their contribution is challenging, as different population synthesis codes yield divergent results in terms of the black hole mass function and merger rates. Furthermore, we characterize the host galaxies where massive black holes form in terms of their dark matter, stellar mass, and metallicity.
Ultimately, we demonstrate that the merger and birth rate densities of binary black holes increase with redshift until z = 5. This cosmic evolution is a crucial signature with significant implications for future detectors such as LISA, the Einstein Telescope, and Cosmic Explorer, which will be capable of probing the high-redshift Universe.
@@ 1 Introduction
The detection of gravitational waves (GWs) in 2015 (Abbott et al. 2016) marked the beginning of a new era in astrophysics, unveiling the mysterious universe of black holes (BHs). Since then, the LIGO-VIRGO-KAGRA (LVK) Collaboration has published nearly 180 binary BHs (BBHs) (Abbott et al. 2023a; The LIGO Scientific Collaboration et al. 2025c), providing crucial insights into the fundamental processes that govern the formation and evolution of these enigmatic compact objects (Abbott et al. 2017, 2019, 2021, 2023b, 2024; The LIGO Scientific Collaboration et al. 2025b). Despite remarkable progress being made, the properties of the merging BBHs detected by LVK, such as their mass and spin distributions, continue to challenge theoretical models. For example, the absence of a high-mass cutoff in the primary BH mass distribution above $\sim 60\,\rm M_{\odot}$ (Farmer et al. 2019, 2020; van Son et al. 2020; Marchant and Moriya 2020; Callister and Farr 2024) has raised questions about the definite existence of a pair-instability mass gap. This gap, which is predicted by stellar evolution models, does not appear in the LVK catalog (Abbott et al. 2023b; The LIGO Scientific Collaboration et al. 2025b), which includes several detections of BHs within the gap region (Abbott et al. 2019). Nevertheless, the precise astrophysical channel responsible for the formation of massive stellar BBHs (MBBHs) with $\rm M_{\rm BH}>50\,\rm M_{\odot}$ remains elusive, with a variety of mechanisms proposed to explain the most massive GW events. These mechanisms include isolated binary evolution (Bethe and Brown 1998; Belczynski et al. 2008; Dominik et al. 2012, 2013, 2015; Belczynski et al. 2016; Schneider et al. 2017; Giacobbo and Mapelli 2018; Marassi et al. 2019; Graziani et al. 2020a; Broekgaarden et al. 2022; Franciolini et al. 2024), direct star collisions (Di Carlo et al. 2019; Renzo et al. 2020; Kremer et al. 2020b, a; Costa et al. 2022; Ballone et al. 2023), three-body interactions (Antonini et al. 2017; Arca Sedda et al. 2021; Vigna-Gómez et al. 2021), and dynamical formation channels involving star clusters (Mapelli 2016; Kumamoto et al. 2020) and the disks of active galactic nuclei (McKernan et al. 2012; Bartos et al. 2017; McKernan et al. 2018; Tagawa et al. 2020b, a; Ishibashi and Gröbner 2020; Vaccaro et al. 2024; McKernan et al. 2025). Among the proposed mechanisms, dynamical formation in dense stellar environments - such as globular clusters (GCs) and nuclear star clusters - has attracted significant interest over the past few decades. These environments are particularly well suited for the formation of hierarchical BH merger chains, where the compact remnants of massive stars can undergo successive coalescences, leading to the formation of increasingly massive BHs (Miller and Hamilton 2002; Holley-Bockelmann et al. 2008; Giersz et al. 2015; Antonini et al. 2016; Antonini and Rasio 2016; Mapelli et al. 2021, 2022; Chattopadhyay et al. 2023; Kritos et al. 2023a; Torniamenti et al. 2024). Indeed, due to mass segregation, BHs gradually migrate toward the cluster core on sub-gigayear timescales, resulting in the formation of a BH-dominated central region (Spitzer 1969; Kulkarni et al. 1993; Sigurdsson and Hernquist 1993). In such BH-dominated cores, dynamically hard BBHs are promptly formed through three-body interactions, and further hardened by GW capture, binary-single, and binary-binary scattering events (Phinney and Sigurdsson 1991; Portegies Zwart and McMillan 2000; Peuten et al. 2016; Arca Sedda et al. 2018; Zocchi et al. 2019; Kremer et al. 2020b; Antonini and Gieles 2020). Additionally, dynamical mechanisms can produce unique features that could help to distinguish the contribution of this formation channel from that of isolated binary evolution, such as large spin magnitudes inherited from the premerger orbital angular momentum (Gerosa and Berti 2017) and non-negligible eccentricities (Nishizawa et al. 2016; Breivik et al. 2016; Nishizawa et al. 2017; Samsing et al. 2019; Martinez et al. 2020; Rom et al. 2024). However, the dynamical formation of BBHs presents several challenges. A major obstacle is the relativistic kick imparted to merger remnants at birth (Campanelli et al. 2007; Lousto and Zlochower 2011), which can exceed the escape velocity of their host clusters. This often results in the ejection of the remnants from their environment, thereby disrupting the hierarchical merger chain (Favata et al. 2004; Holley-Bockelmann et al. 2008; Kesden et al. 2010). The interplay between the properties of the host environment and the strength of the kicks plays a critical role in shaping the mass distribution of the BH population resulting from this formation pathway (Islam and Wadekar 2025). Unfortunately, due to the large distances involved and limited sky localization accuracy, confidently determining the orbital properties and specific birth environments of BBHs remains challenging. This uncertainty hampers our ability to fully understand the true origin of MBBHs and to rule out alternative formation channels, such as the binary evolution of Population III stars (Kinugawa et al. 2014, 2016; Tanikawa et al. 2021; Wang et al. 2022; Tanikawa et al. 2022; Costa et al. 2023; Santoliquido et al. 2023; Mestichelli et al. 2024; Liu et al. 2024) and primordial BHs (De Luca et al. 2021; Ng et al. 2022; Chen and Hall 2024; De Luca et al. 2025) as contributors to the observed GW events. As the number of detected mergers continues to grow with ongoing and future runs, the prospect of disentangling the contributions from different formation mechanisms becomes increasingly attainable. Bayesian inference methods (Thrane and Talbot 2019; Abbott et al. 2023b; Franciolini and Pani 2022; Franciolini et al. 2024) and population synthesis models (Hurley et al. 2000; Portegies Zwart and Verbunt 1996; Mapelli et al. 2013; Spera et al. 2015; Giacobbo et al. 2018; Elbert et al. 2018; Breivik et al. 2020; Riley et al. 2022; Iorio et al. 2023) enable the inference of the intrinsic properties of observed BBHs by comparing observational data with predictions derived from various astrophysical scenarios and binary initial conditions. These statistical comparisons have opened new avenues for uncovering the true formation history behind these mergers, despite the significant uncertainties associated with stellar and binary evolution (Iorio et al. 2023). In this work, we focus on the role of the dynamical formation scenario in dense stellar environments, with particular emphasis on the contribution of BBHs formed in GCs to the observed population of BH mergers. Additionally, we investigate how the distinctive properties of BHs resulting from hierarchical mergers may provide new insights into the astrophysical processes underlying the formation of MBBHs. To investigate this, various approaches have been employed. Traditionally, the numerical complexity of direct $N$-body and Monte Carlo simulations has limited studies of cluster dynamics to a relatively small number of cases (Spurzem 1999; Rodriguez et al. 2015, 2016a, 2016b; Wang et al. 2015, 2016; Chatterjee et al. 2017; Askar et al. 2017; Rodriguez et al. 2019; Banerjee et al. 2020; Arca Sedda et al. 2023, 2024). However, the recent development of fast semi-analytical codes, also known as cluster population synthesis (CPS) codes – such as B-POP (Arca Sedda et al. 2021), cBHBd (Antonini and Gieles 2020; Antonini et al. 2023), FASTCLUSTER (Mapelli et al. 2021), QLUSTER (Gerosa 2023), and RAPSTER (Kritos et al. 2024) – has made it possible to conduct cluster population studies within a reasonable time-frame. By combining CPS predictions with models of the cosmic cluster formation rate (Mapelli et al. 2022; González Prieto et al. 2022; Kritos et al. 2024; Torniamenti et al. 2024; Mestichelli et al. 2024), one can rapidly estimate the average merger rate density of dynamically formed BBHs as a function of redshift. However, this approach does not provide detailed information about the spatial distribution or the chemical properties of the galaxies hosting globular clusters. Understanding the properties of host galaxies is crucial for improving our knowledge of GCs and, by extension, of MBBH formation (see Krumholz et al. 2019, for a review). Several key aspects of cluster formation are known to vary with the galactic environment, including the cluster formation efficiency (Ginsburg and Kruijssen 2018), the potential high-mass truncation of the cluster initial mass function (Wainer et al. 2022), and the mass–radius relation (Brown and Gnedin 2021). Moreover, the time evolution of individual star clusters is strongly influenced by a range of environmental factors (e.g., Rossi et al. 2016; Suin et al. 2022). Therefore, a realistic model of GC formation and evolution must incorporate the evolving, galaxy-specific environmental conditions, and do so across a statistically significant sample of galaxies to adequately represent their diversity. By incorporating the progression of chemical enrichment in galaxies throughout their assembly, hydrodynamic galaxy formation simulations offer a powerful tool to study the population of dynamically formed BBHs emerging from dense stellar environments. However, due to their high computational cost, these simulations require a trade-off between mass and spatial resolution, the size of the simulated volume, and the redshift range covered. Currently, most high-resolution zoom-in simulations – those with sub-parsec resolution capable of capturing star cluster formation – are limited to small cosmological volumes and/or high redshifts (e.g., Boley et al. 2009; Kimm et al. 2016; Lahén et al. 2019; Calura et al. 2022, 2025; Sugimura et al. 2024). When the mass resolution is insufficient to resolve star cluster formation within galaxies, semi-analytical models become essential to complement these simulations. Studies utilizing this approach have already been carried out, demonstrating promising agreement with the observed GC populations in nearby galaxies (e.g., Li and Gnedin 2014; Pfeffer et al. 2018). This study aims to investigate the physical properties of dynamically formed BBHs originating in GCs during the assembly of the Local Group (LG) simulated by the galaxy formation model GAMESH (Graziani et al. 2017). By estimating their merger rate density and examining their birth environments, we quantify the contribution of this dynamical formation channel to the GW events detected by the LVK Collaboration. Additionally, our analysis enables the identification of the host galaxies of these BBHs and the characterization of their physical properties. To this end, in Section 2.1 we outline the adopted galaxy formation model, in Section 2.2 we detail the coupling between the CPS framework and the galaxy evolution simulation, and in Section 2.3 we describe our cluster evolution models. Our main findings are presented in Section 3, followed by an extended comparison with previous studies in Section 4. This discussion is especially timely because of the recent release of O4 data (Abac and others 2025).
@@ 2 Methods
In this study, we adopt a LG-like volume simulation, named GAMESH (Graziani et al. 2015) (Section 2.1) and propose a novel semi-analytic coupling approach for GC formation (Section 2.2), enabling the placement of cluster populations within each galaxy. Finally, we integrate the time evolution of GCs using the RAPSTER and FASTCLUSTER codes (Section 2.3), marking their first application within the context of a numerical galaxy formation simulation.